OpenStudy (tal_tal):

Part A: Michael bought vegetables that weighs 4 and 1 over 8 pounds. How many ounces does the vegetables weigh? Show your work. (5 points) [16 ounces = 1 pound] Part B: A running tap dispenses 0.16 gallons of water every second. How many pints of water is dispensed after 25 seconds? Show your work. (5 points) [1 gallon = 4 quarts, 1 quart = 2 pints]

OpenStudy (tal_tal):

@bibby I'm sorry last one for the day

OpenStudy (mathmale):

So, you start with vegetables that all together weigh \[4\frac{ 1 }{ 8 }lb.\]

OpenStudy (mathmale):

Write the appropriate conversion factor (pounds to ounces). Multiply your \[4\frac{ 1 }{ 8 }lb\]

OpenStudy (mathmale):

by that conversion factor. You'll know you're on the right track if "lb" disappears and "oz" shows up in its place.

OpenStudy (tal_tal):

Wait which one is part B

OpenStudy (mathmale):

Go back to your original posting. It was you who presented Part B.

OpenStudy (tal_tal):

Okay I got it thank you so much!!!

OpenStudy (tal_tal):

Wait @bibby sorry to bother you but can you help me with part B

OpenStudy (bibby):

find the amount of gallons in 25 seconds by multiplying: \(\dfrac{0.16gal}{1\cancel{sec}}*25 \cancel{sec}\) Then convert that much gallons into pints

OpenStudy (tal_tal):

Thank you